OpenStudy (anonymous):

Solve the triangle. Round lengths to the nearest tenth and angle measures to the nearest degree. a = 6, c = 11, B = 109° A. b = 14.1, A = 24°, C = 47° B. b = 19.9, A = 22°, C = 49° C. b = 17, A = 26°, C = 45° D. no triangle

OpenStudy (anonymous):

@amistre64

OpenStudy (anonymous):

@Abhisar

OpenStudy (anonymous):

@aaronq

OpenStudy (anonymous):

@agent0smith

OpenStudy (anonymous):

@dan815

OpenStudy (anonymous):

@ajprincess @precal

OpenStudy (anonymous):

I know it is not C.

OpenStudy (anonymous):

I think it is B. what do you think @ajprincess

OpenStudy (anonymous):

Use the law of cosines and law of sines to figure this out. Law of cosines \[b ^{2}=a ^{2}+c ^{2}-2ac \cos B\] Law of sines\[\frac{ \sin A }{ a}=\frac{ \sin B }{ b }=\frac{ \sin C }{ c }\]

OpenStudy (anonymous):

Well its either A or C

OpenStudy (anonymous):

It can't be both. Plug the numbers into the law of cosine equation.